OpenStudy (anonymous):

Express in terms of i. 2√-36

OpenStudy (anonymous):

12i

OpenStudy (anonymous):

can you plz explain the steps you took to get that

OpenStudy (anonymous):

\[2\sqrt{36i^2}=2*6i=12i\]

OpenStudy (anonymous):

36 is a perfect square, 6 is its square root

OpenStudy (anonymous):

√-1 = i , √36 =6 ->

OpenStudy (anonymous):

square root of i^2=i

OpenStudy (anonymous):

√-1√36 = √-36=6i

OpenStudy (anonymous):

thank u